What is the meaning of versin(x)?
I found a question while solving Trigonometry by SL Loney.

Grade:12

2 Answers

ng29
9 years ago
http://commons.wikimedia.org/wiki/File:Circle-trig6.svg
ng29
9 years ago
i have send u alink check it out and ur doubt will be cleared 
Versine: versin(θ)=1-cos(θ)
Vercosine: vercosin(θ)=1+cos(θ)
Coversine: coversin(θ)=1-sin(θ)
Covercosine: covercosine(θ)=1+sin(θ)
Haversine: haversin(θ)=versin(θ)/2
 
 
plz approve if useful
plz approve if useful